OpenStudy (anonymous):

Hi, can someone help me with parallel vectors? I have vector PR=<-1,10>, and I want to find another vector parallell to it.

ganeshie8 (ganeshie8):

multiply it wid a scalar number

OpenStudy (anonymous):

so would <-2,20> be a good answer?

ganeshie8 (ganeshie8):

yup ! in general, say \(\overrightarrow{PR}\) is a vector, then the vector \(k\overrightarrow{PR}\) will be parallel vector. \(k\) can be any scalar. if k is positive, then the resulting vector will be in same direction, if k is negative, then the resulting vector will be in opposite direction.

OpenStudy (anonymous):

And also the dot product should equal 1 right? To show the angle between 2 vectors in 0 deg

ganeshie8 (ganeshie8):

oh to show the angle between them is 0, u need to show \(\cos(\theta) = \cos(0) = 1\)

ganeshie8 (ganeshie8):

u may use dot product formula for that

ganeshie8 (ganeshie8):

\(\vec{A} . \vec{B} = |\vec{A}| |\vec{B}|\cos(\theta)\) plugin values of both vectors

ganeshie8 (ganeshie8):

\(<-1, 10> . <-2, 20> = \sqrt{(-1)^2 + 10^2}\sqrt{(-2)^2+20^2} \cos(\theta)\) simplify and solve \(\cos(\theta)\)

OpenStudy (anonymous):

Thank you! (;

ganeshie8 (ganeshie8):

np :) u knw how to do the left hand side right ?

OpenStudy (anonymous):

Yes.

ganeshie8 (ganeshie8):

good :)